OpenStudy (anonymous):

Show your work to prove that the equation y = 12x + 8 correctly represents the function for the data in this table.

OpenStudy (anonymous):

Post the table that the question is referring to and I can help.

OpenStudy (anonymous):

The easiest way to prove that the equation represents the data is to plug the "x" value in. y = 12(0) + 8 = 0 + 8 = 8 This confirms the first line where x = 0 and y = 8 Do you understand? If so, do the same thing for the other problems, and I can confirm if you are right or not.